OpenStudy (anonymous):

A train leaves Lexington for Indianapolis, 200 miles away, at 3:00 PM and averages 60 mph. A second train traveling on an adjacent track leaves Indianapolis for Lexington at 5:30 PM and averages 40 mph. At what time will the trains meet?

OpenStudy (anonymous):

Set this up. Let t= time after 3:00 (we will assume the two locations are in the same time zone). So the distances traveled will have to total 200 miles. That means\[60t+40(t-5/2)=200\]which I think you should be able to solve.

OpenStudy (anonymous):

If my mental math is any good, t=3, so the trains meet at 6:00 PM.

OpenStudy (anonymous):

Where did the -5/2 come from?

OpenStudy (anonymous):

That's the two and a half hour head start the first train gets. The second train doesn't start until 5:30, so its time traveled is t-2.5=t-5/2.

OpenStudy (anonymous):

Ohh okay, I understand now! Setting up these type of problems are a bit difficult

OpenStudy (anonymous):

Just takes some practice. Keep after it!

OpenStudy (anonymous):

Thank you!